# Explainable Concept Generation through Vision-Language Preference Learning

### v2

@@ Abstract
Concept-based explanations have become a popular choice for explaining deep neural networks post-hoc because, unlike most other explainable AI techniques, they can be used to test high-level visual “concepts” that are not directly related to feature attributes. For instance, the concept of “stripes” is important to classify an image as a zebra. Concept-based explanation methods, however, require practitioners to guess and collect multiple candidate concept image sets, which can often be imprecise and labor-intensive. Addressing this limitation, in this paper, we frame concept image set creation as an image generation problem. However, since naively using a generative model does not result in meaningful concepts, we devise a reinforcement learning-based preference optimization (RLPO) algorithm that fine-tunes the vision-language generative model from approximate textual descriptions of concepts. Through a series of experiments, we demonstrate the capability of our method to articulate complex and abstract concepts which aligns with the test class that are otherwise challenging to craft manually. In addition to showing the efficacy and reliability of our method, we show how our method can be used as a diagnostic tool for analyzing neural networks.
@@ 1 Introduction
In an era where black box deep neural networks (DNNs) are becoming seemingly capable of performing general enough tasks, our ability to explain their decisions post-hoc has become even more important before deploying them in the real world. Among many use cases, explaining the behavior of DNNs enable engineers and regulatory bodies to assess the correctness of a DNN’s decision-making process and characterize the limits of what the DNN knows. Thus, an explainable AI (XAI) method’s ability to correctly communicate information in a human-centric way is essential for its usefulness.
Humans utilize high-level concepts as a medium for providing and perceiving explanations. In this light, post-hoc concept-based explanation techniques, such as Testing with Concept Activation Vectors (TCAV) (Kim et al., 2018), have gained great popularity in recent years. Their ability to use abstractions that are not necessarily feature attributes or some pixels in test images helps with communicating these high-level concepts with humans. For instance, as demonstrated in TCAV, the concept of stripes is important to explain why an image is classified as a zebra, whereas the concept of spots is important to explain why an image is classified as a jaguar. Given 1) a set of such high-level concepts, represented as sample images (e.g., a collection of stripe images and a collection of spot images) and 2) test images of the class (e.g., zebra images), TCAV assigns a score to each concept on how well the concept explains the class decision (i.e., zebra).
Figure 1: (a) Our proposed algorithm, RLPO, iteratively refines the concepts $c_{i}$ that can be generated by a Stable Diffusion (SD) model by optimizing SD weights based on an action $a_{i}$. Each step in this update process provides an explanation at a different level of abstraction. (b) Three concepts identified by our approach for the zebra class. Concepts are represented as images generated by SD.
@@
Although concept-based XAI methods are a good representation, their requirement to create collections of candidate concept sets necessitate the human to know which concepts to test for. This is typically done by guessing what concepts might matter and manually extracting such candidate concept tests from existing datasets. While the stripe-zebra analogy is attractive as an example, where it is obvious that stripes is important to predict zebras, in most applications, we cannot guess what concepts to test for, limiting the usefulness of concept-based methods in testing real-world systems. Additionally, even if a human can guess a few concepts, it does not encompass most concepts a DNN has learned because the DNN was trained without any human intervention. Therefore, it is important to automatically find human-centric concepts that matter to the DNN’s decision-making process.
Figure 2: We can generate concepts that encompass both human-defined and retrieved concepts. Note that retrieved concepts are very similar to class images, making them less useful as concepts.
@@
As attempts to automatically discover and create such concept sets, several work has focused on segmenting the image and using these segments as potential concepts, either directly (Ghorbani et al., 2019) or through factor analysis (Fel et al., 2023; 2024). In such methods, which we refer to as retrieval methods, because the extracted concept set is already part of the test images as shown in Fig. 2 (Retrieved concepts), it is difficult for them to imagine concepts that do not have a direct pixel-level resemblance to the original image class. For instance, it is more likely that such methods provide patches of zebra as concepts instead of stripes.
@@
By departing from existing concept set creation practices of human handcrafting and retrieval, we redefine concept set creation as a concept generation problem. Modern generative models such as stable diffusion (SD) can be used for generating realistic images. Nevertheless, since a generative mode generates arbitrary images, we need to carefully guide it to generate explanatory images. One obvious approach is to engineer long, descriptive text prompts to generate concepts. However, engineering such prompts is not realistic. Therefore, to automate this process, as shown in Fig. 1, we propose a method, named reinforcement learning-based preference optimization (RLPO), to update SD weights. At its core, we devised a deep reinforcement learning algorithm gradually update SD weights to generate concept images that have a higher explanation score. The contributions of this paper can be summarized as follows:
@@
    We propose a method, named RLPO, to “generate” concepts that truly matters to the neural network. Some of these concepts would not pop up in humans head until they see them. Also, unlike existing retrieval methods, RLPO can generate concepts that are not part of test images (Fig. 2).
@@
    Because of how we use model’s preference to gradually make the SD process closer to the target class from a high-level concept, for the first time in concept-based XAI, we can generate concepts with different abstraction levels.
@@
    In addition to demonstrating the novelty, abstractness, and diversity of generated concepts, we experimentally verify their generalizability using an NLP sentiment analysis tasks and the actionability by leveraging concepts as a tool for fine-tuning.
@@ 2 Preliminaries and Related Work
Testing with Concept Activation Vectors (TCAV): The TCAV score quantifies the importance of a “concept” for a specific class in a DNN classifier (Kim et al., 2018). Here, a concept is defined broadly as a high-level, human-interpretable idea such as stripes, sad faces, etc. A concept (e.g., stripes), $c$, is represented by sample images, $X_{c}$ (e.g., images of stripes). In TCAV, a human has manually collected these sample concept images based on educated guesses, whereas our objective is to automatically generate them. For a given set of test images, $X_{m}$ (e.g., zebra images), that belongs to the same decision class (e.g., zebra), $m$, TCAV is defined as the fraction of test images for which the model’s prediction increases in the “direction of the concept.” By decomposing the DNN under test as $f(x)=f_{2}(f_{1}(x))$, where $f_{1}(x)$ is the activation at layer $l$, TCAV score is computed as,
$TS_{c,m}=\frac{1}{|X_{m}|}\sum_{X_{m}}\mathbb{I}\left(\frac{\partial\text{output}}{\partial\text{activations}}\cdot{(c\enspace\text{direction})}>0\right)=\frac{1}{|X_{m}|}\sum_{x_{i}\in X_{m}}\mathbb{I}\left(\frac{\partial f(x_{i})}{\partial f_{1}(x_{i})}\cdot v>0\right)$ (1)
Here, $\mathbb{I}$ is the indicator function that counts how often the directional derivative is positive. Concept activations vector (CAV), $v$, is the normal vector to the hyperplane that separates activations of concept images, $\{f_{1}(x);x\in X_{c}\}$, from activations of random images, $\{f_{1}(x);x\in X_{r}\}$. Refer to Appendix C.1 for details on the TCAV settings.
ACE (Ghorbani et al., 2019) introduced a way to automatically find concepts by extracting relevant concepts from the input class. It used segmentation over different resolution to get a pool of segments and then grouped them based on similarity to compute TCAV scores. Though the ACE concepts are human understandable, they are noisy because of the segmentation and clustering errors. As a different method, EAC (Sun et al., 2024) extracts concepts through segmentation. CRAFT (Fel et al., 2023) introduced a recursive strategy to detect and decompose concepts across layers. Lens (Fel et al., 2024) elegantly unified concept extraction and importance estimation as a dictionary learning problem. However, since all these methods obtain concepts from test images, the concepts they generate tend to be very similar to the actual class (e.g., a patch of zebra as a concept to explain the zebra instead of stripes as a concept), making it challenging to maintain the “high-level abstractness” of concepts. In contrast, we generate concepts from a generative model. Under generative models, LCDA (Yan et al., 2023) simply queries an LLM to get attributes but does not generate concepts.
@@
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\left(r(s_{t},a_{t})+\gamma\max_{a^{\prime}}Q_{\text{target}}(s_{t+1},a^{\prime})-Q(s_{t},a_{t})\right).$ (2)
Here, $s_{t}$ is the state at step $t$, $a_{t}$ is the action taken in state $s_{t}$, and $r_{t}$ is the reward received after taking action $a_{t}$. The parameters $\alpha$ and $\gamma$ are learning rate and discount factor, respectively.
Preference Optimization: Directly optimizing generative models with preference data was first introduced in Direct Preference Optimization (DPO) (Rafailov et al., 2024). It is a technique used to ensure models, such as large language models, learn to align its outputs with human preference by asking a human which of its generated output is preferred. This technique was later extended to diffusion models in Diffusion-DPO (Wallace et al., 2023), where they updated Stable Diffusion XL model using Pick-a-Pic dataset (human preferred generated image dataset). Unlike traditional image or text generation tasks, where the dataset for human preferred outputs are readily available, it is hard to have a general enough dataset for XAI tasks. To counter this problem, we provide preference information by using the TCAV score instead of a human, and use it to align the text-to-image generative model to generate concept images that matters for the neural network under test.
@@ 3 Methodology: Reinforcement Learning-based Preference Optimization
Our objective is to find a set of concept images, $\mathcal{C}$, that maximize the TCAV scores, $TS_{c,m}$. To this end, we leverage the state-of-the-art text-to-image generative models to generate high quality explainable concepts. However, because the search space of potential text prompts is too large, we use deep RL to guide the image generation process. As described in Fig. 3 and Algorithm 1, our algorithm, RLPO, picks potential keywords from an automatically generated list of keywords using RL and optimizes stable diffusion weights to generate images that have a preference for higher TCAV scores. This process is described in the rest of this section.
Algorithm 1 The RLPO algorithm. Appendix C.2 for the expanded algorithm.
1:Input: Set of test images, $f(\cdot)$
2:Run pre-processing and get the seed prompts (action space)
3:for each episode do
4:     for each time step $t$ do
5:         Execute $a_{t}$ and pick a keyword
6:         Generate image groups $G_{1}$ $G_{2}$
7:         Evaluate TCAV scores $TS_{1}$ $TS_{2}$
8:         Update SD based on better score
9:         Compute reward
10:     end for
11:end for
12:Output: Set of concept images
Notation: Our framework contains three core deep learning models: the network under test $f(\cdot)$, the image generator $g(\cdot)$, and the deep RL network $h(\cdot)$. First, we have a pre-trained neural network classifier that we want to explain. We then have a generative neural network, whose purpose is generating concept image sets, given some text prompts. In this paper, we use Stable Diffusion (SD) v1-5 as the generator as it is a state-of-the-art generative model that can generate realistic images. If the weights of the SD model are $w$, for a small constant $\lambda$, we augment it as $w+\lambda ab$, where $A$ and $B$ are low-rank matrices that we fine-tune using preference optimization (Rafailov et al., 2024). The core search algorithm that we train is a DQN.
@@ 3.1 The Rationales Behind Design Choices
Before presenting the algorithm in detail, we provide the rationale for design choices, which are validated through ablation studies in Section 4.1-4.2, and comparisons in the Section 4.3-4.4.
Rationale 1: Why concept generation is a better idea. Let us denote the set of human-interpretable concepts that the $f(\cdot)$ has learned be $\mathcal{C}_{N}$. If we use concept-based explanation the traditional way (Kim et al., 2018; Schut et al., 2023), then the end users need to manually guess what concepts to test for. Automatically retrieving the concept set by segmenting test images (Sun et al., 2024) also results in a limited concept set. In contrast, a SOTA generative model can generate high quality images. We provide more theoretical insights in Appendix B.
Rationale 2: Why a deep RL-controlled VLM fine-tuning for generating concepts is a better idea. “A picture is worth a thousand words but words flow easier than paint.”
As the saying goes, “a picture is worth a thousand words,” it is much easier for people to explain and understand high-level concepts when images are used instead of language. For instance, we need a long textual description such as “The circles are centered around a common point, with alternating red and white colors creating a pattern” to describe a simple image of a dart board (i.e., Target Co. logo). Therefore, we keep our ultimate concept representation as images. However, controlling a generative model from visual inputs is much harder. However, since human language can be used as a directed and easier way to seed our thought process, as the saying goes, “words flow easier than paint,” we control the use of text prompts. Since the vastness of the search space cannot be handled by most traditional search strategies, we resort to a DQN for controlling text. Since simple text alone cannot generate complex, high-level visual concepts, in each DQN update step, we use preference optimization to further guide the search process towards more preferred outcome, allowing the DQN to focus on states similar to the target. This approach improves our starting points for each DQN episode, enabling more efficient search and incremental progress towards the desired target.

### v3

@@ Abstract
Understanding the inner representation of a neural network helps users improve models. Concept-based methods have become a popular choice for explaining deep neural networks post-hoc because, unlike most other explainable AI techniques, they can be used to test high-level visual “concepts” that are not directly related to feature attributes. For instance, the concept of “stripes” is important to classify an image as a zebra. Concept-based explanation methods, however, require practitioners to guess and manually collect multiple candidate concept image sets, making the process labor-intensive and prone to overlooking important concepts. Addressing this limitation, in this paper, we frame concept image set creation as an image generation problem. However, since naively using a standard generative model does not result in meaningful concepts, we devise a reinforcement learning-based preference optimization (RLPO) algorithm that fine-tunes a vision-language generative model from approximate textual descriptions of concepts. Through a series of experiments, we demonstrate our method’s ability to efficiently and reliably articulate diverse concepts that are otherwise challenging to craft manually. Github: https://github.com/aditya-taparia/RLPO
Machine Learning, ICML, Concept based Explainable AI, TCAV, Vision-Language Models, Reinforcement Learning, Preference Optimization
@@ 1 Introduction
In an era where black box deep neural networks (DNNs) are becoming seemingly capable of performing complex tasks, our ability to understand their internal representations post-hoc has become even more important before deploying them in the real world. Among many use cases, such revelations help engineers in further improving models and regulatory bodies in assessing their correctness. To help these users, developing methods that communicate information in a human-centric way is essential for ensuring usefulness.
@@
Humans utilize high-level concepts as a medium for providing and perceiving explanations. In this light, post-hoc concept-based explanation techniques, such as Testing with Concept Activation Vectors (TCAV) (Kim et al., 2018), have gained great popularity in recent years. Their ability to use abstractions that are not necessarily feature attributes or some pixels in test images helps with communicating these high-level concepts with humans. For instance, as demonstrated in TCAV, the concept of stripes is important to explain why an image is classified as a zebra, whereas the concept of spots is important to explain why an image is classified as a jaguar. Given 1) a set of such high-level concepts, represented as sample images (e.g., a collection of stripe and spot images) and 2) test images of the class (e.g., zebra images), TCAV assigns a score to each concept on how well the concept explains the class decision (i.e., zebra). Although concept-based explanations are a good representation, their requirement to create collections of candidate concept sets necessitate the human to know which concepts to test for. This is typically done by guessing what concepts might matter and manually extracting such candidate concept tests from existing datasets. While the stripe-zebra analogy is attractive as an example, where it is obvious that stripes is important to predict zebras, in most applications, we cannot guess what concepts to test for, limiting the usefulness of concept-based methods in testing real-world systems. Additionally, even if a human can guess a few concepts, it does not encompass most concepts a DNN has learned because the DNN was trained without any human intervention. Therefore, it is important to automatically find concepts that matter to the DNN’s decision-making process.
Figure 1: Humans can imagine (green) a few concepts to understand neural networks’ representations (purple). Some other concepts can be retrieved from test images themselves through, for instance, segmentation (orange). However, if we generate concepts (blue), they will capture even a broader set of concepts. RLPO is designed to make this generation process more targeted toward neural networks’ representations (purple $\cap$ blue).
Figure 2: (a) Our proposed algorithm, RLPO, iteratively refines the concepts $c_{i}$ that are generated by a Stable Diffusion (SD) model by optimizing SD weights based on an action $a_{i}$. Each step in this update process provides an explanation at a different level of abstraction (Appendix C.4). Please refer to the supplementary video for better understanding. (b) Three concepts identified by our approach for the zebra class. Concepts are represented as images generated by SD.
@@
As attempts to automatically discover and create such concept sets, several works has focused on segmenting the image and using these segments as potential concepts, either directly (Ghorbani et al., 2019) or through factor analysis (Fel et al., 2023, 2024). In such methods, which we refer to as retrieval methods, because the extracted concept set is already part of the test images as shown in Fig. 1 (Retrieved concepts), it is difficult for them to imagine concepts that do not have a direct pixel-level resemblance to the original image class. For instance, it is more likely that such methods provide patches of zebra as concepts instead of stripes.
@@
By departing from existing concept set creation practices of human handcrafting and retrieval, we redefine concept set creation as a concept generation problem. Modern generative models such as stable diffusion (SD) can be used to generate realistic images. Nevertheless, since a generative model generates arbitrary images, we need to carefully guide the SD to generate explanatory images. One obvious approach is to engineer long, descriptive text prompts to generate concepts. However, engineering such prompts is not realistic. Therefore, to automate this process, as shown in Fig. 2, we propose a method, named reinforcement learning-based preference optimization (RLPO), to guide SD model. At its core, we devised a deep reinforcement learning algorithm, which gradually update SD weights to generate concept images that have a higher explanation score. The contributions of this paper can be summarized as follows:
@@
    We propose a method, named RLPO, to “generate” concepts that truly matter to the neural network (refer to the video on github). These include concepts that are challenging for humans or retrieval methods to anticipate (Fig. 1).
@@
    We qualitatively (computational metrics and human surveys) and quantitatively demonstrate that RLPO can reveal networks’ internal representations.
@@
    Additionally, as use cases, we show how generated concepts can be utilized to improve models, understand model bias, and also demonstrate the generalizability of RLPO on tasks such as NLP sentiment analysis.
Notably, due to the modularity of our approach, any component in our framework (e.g., SD) can be seamlessly replaced with an improved version (e.g., SD3 (Esser et al., 2024) or Flash Diffusion (Chadebec et al., 2025)).
@@ 2 Preliminaries and Related Work
Testing with Concept Activation Vectors (TCAV): The TCAV score quantifies the importance of a “concept” for a specific class in a DNN classifier (Kim et al., 2018). Here, a concept is defined broadly as a high-level, human-interpretable idea such as stripes, sad faces, etc. A concept (e.g., stripes), $c$, is represented by sample images, $X_{c}$ (e.g., images of stripes). In TCAV, a human has manually collected these sample concept images based on educated guesses, whereas our objective is to automatically generate them. For a given set of test images, $X_{m}$ (e.g., zebra images), that belong to the same decision class (e.g., zebra), $m$, TCAV is defined as the fraction of test images for which the model’s prediction increases in the “direction of the concept.” By decomposing the DNN under test as $f(x)=f_{2}(f_{1}(x))$, where $f_{1}(x)$ is the activation at layer $l$, TCAV score is computed as,
$\begin{split}TS_{c,m}&=\frac{1}{|X_{m}|}\sum_{X_{m}}\mathbb{I}\left(\frac{\partial\text{output}}{\partial\text{activations}}\cdot{(c\enspace\text{direction})}>0\right)\\ &=\frac{1}{|X_{m}|}\sum_{x_{i}\in X_{m}}\mathbb{I}\left(\frac{\partial f(x_{i})}{\partial f_{1}(x_{i})}\cdot v>0\right)\end{split}$ (1)
Here, $\mathbb{I}$ is the indicator function that counts how often the directional derivative is positive. Concept activations vector (CAV), $v$, is the normal vector to the hyperplane that separates activations of concept images, $\{f_{1}(x);x\in X_{c}\}$, from activations of random images, $\{f_{1}(x);x\in X_{r}\}$. Refer to Appendix B.1 for details on the TCAV settings.
ACE (Ghorbani et al., 2019) introduced a way to automatically find concepts by extracting relevant concepts from the input class. It used segmentation over different resolution to get a pool of segments and then grouped them based on similarity to compute TCAV scores. Though the ACE concepts are human understandable, they are noisy because of the segmentation and clustering errors. As a different method, EAC (Sun et al., 2024) extracts concepts through segmentation. CRAFT (Fel et al., 2023) introduced a recursive strategy to detect and decompose concepts across layers. Lens (Fel et al., 2024) elegantly unified concept extraction and importance estimation as a dictionary learning problem. However, since all these methods obtain concepts from class images, the concepts they generate tend to be very similar to the actual class (e.g., a patch of zebra as a concept to explain the zebra, instead of stripes as a concept), making it challenging to maintain the “high-level abstractness” of concepts. In contrast, we generate concepts from a generative model.
@@
$\begin{split}Q(s_{t},a_{t})&\leftarrow Q(s_{t},a_{t})+\alpha\Big{(}r(s_{t},a_{t})\\ &+\gamma\max_{a^{\prime}}Q_{\text{target}}(s_{t+1},a^{\prime})-Q(s_{t},a_{t})\Big{)}\end{split}$ (2)
Here, $s_{t}$ is the state at step $t$, $a_{t}$ is the action taken, and $r_{t}$ is the reward obtained after taking action $a_{t}$. The parameters $\alpha$ and $\gamma$ are learning rate and discount factor, respectively. DQNs are used for controlling robots (Tang et al., 2025; Senanayake, 2024; Charpentier et al., 2022), detecting failures (Sagar et al., 2024), etc.
Preference Optimization: Optimizing generative models with preference data was first introduced in Direct Preference Optimization (DPO) (Rafailov et al., 2024). It is a technique used to ensure models, such as large language models, learn to align its outputs with human preference by asking a human which of its generated output is preferred. This technique was later extended to diffusion models in Diffusion-DPO (Wallace et al., 2024), where they updated Stable Diffusion XL model using Pick-a-Pic dataset (human preferred generated image dataset). Unlike traditional image or text generation tasks, where the dataset for human preferred outputs are readily available, it is hard to have a general enough dataset for XAI tasks. To counter this problem, we provide preference information by using the TCAV score instead of a human, and use it to align the text-to-image generative model to generate concept images that matters for the neural network under test.
Use of VLM in Explanation: Recent advancements in Vision-Language models (VLMs) have open the doors for the use of VLMs in multiple domains, mainly because of their ability to generalize over large amount of data, they can be leveraged to obtain useful information. Work by (Sun et al., 2024) present a novel method combining the Segment Anything Model (SAM) with concept-based explanations, called Explain Any Concept (EAC). This method uses SAM for precise instance segmentation to automatically extract concept sets from images, then it employs a lightweight surrogate model to efficiently explain decision made by any neural network based on extracted concepts. Another work by Yan et al. (Yan et al., 2023) introduced Learning Concise and Descriptive Attributes (LCDA), which leverages Large Language Models (LLMs) to query a set of attributes describing each class and then use that information with vision-language models to classify images. They highlight in their paper that with a concise set of attributes, they can improve the classifier’s performance and also increase interpretability and interactivity for end user.
@@ 3 Methodology: Reinforcement Learning-based Preference Optimization
Our objective is to find a set of concept images, $\mathcal{C}$, that maximize the TCAV scores, $TS_{c,m}$, indicating that the concepts are relevant to the neural networks’ decision-making process. We leverage state-of-the-art text-to-image generative models to generate high-quality explainable concepts. However, because the search space of potential text prompts is too large, we use deep RL to guide the image generation process. As described in Fig. 3 and Algorithm 1, our algorithm, RLPO, uses RL to pick potential keywords from a predefined list and iteratively optimizes stable diffusion weights to generate images that have a preference for higher TCAV scores. This process is described below.
Algorithm 1 The RLPO algorithm. Appendix B.2 for the expanded algorithm.
  Input: Set of test images $f(\cdot)$
  Run pre-processing and obtain seed prompts (action space)
  for each episode do
     for each time step $t$ do
        Execute $a_{t}$ by picking a seed prompt
        Generate image groups $G_{1}$ and $G_{2}$
        Evaluate TCAV scores $TS_{1}$ and $TS_{2}$
        Update SD based on the better score
        Compute reward
     end for
  end for
  Output: Set of concept images
Notation: Our framework contains three core deep learning models: the network under test $f(\cdot)$, the image generator $g(\cdot)$, and the deep RL network $h(\cdot)$. First, we have a pre-trained neural network classifier that we want to explain. We then have a generative neural network, whose purpose is generating concept image sets, given some text prompts. In this paper, we use Stable Diffusion (SD) v1-5 as the generator as it is a state-of-the-art generative model that can generate realistic images. The core search algorithm that we train is a DQN.
@@ 3.1 The Rationales Behind Design Choices
We explain the design choices, which are validated through experiments in Section 4.1-4.4.
Rationale 1: Why concept generation is a better idea. If we use concept-based explanation the traditional way (Kim et al., 2018; Schut et al., 2025), then the end users need to manually guess what concepts to test for. Automatically retrieving the concept set by segmenting test images (Sun et al., 2024) also results in a limited concept set. In contrast, a SOTA generative model can generate high quality images. We provide more theoretical insights in Appendix A.
Rationale 2: Why a deep RL-controlled VLM fine-tuning for generating concepts is a better idea. “A picture is worth a thousand words but words flow easier than paint.” As the saying goes, “a picture is worth a thousand words,” it is much easier for people to explain and understand high-level concepts when images are used instead of language. For instance, we need a long textual description such as “The circles are centered around a common point, with alternating red and white colors creating a pattern” to describe a simple image of a dart board (i.e., Target Co. logo). Therefore, we keep our ultimate concept representation as images. However, controlling a generative model from visual inputs is much harder. Since human language can be used as a directed and easier way to seed our thought process, as the saying goes, “words flow easier than paint,” we control the outcome by using text prompts. Since the vastness of the search space cannot be handled by most traditional search strategies, we resort to a DQN for controlling text. Since simple text alone cannot generate complex, high-level visual concepts, in each DQN update step, we use preference optimization to further guide the search process towards a more preferred outcome, allowing the DQN to focus on states similar to the target. This approach improves our starting points for each DQN episode, enabling more efficient search and incremental progress towards the desired target.